Week of 180430

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Kate (WLCG, DB), Borja (monit), Alberto (monit), Belinda (storage), Jesus (storage)
  • remote: Sang Un (KISTI), Luca T (LHCb), David B (IN2P3), Andrew (NL-T1), Dmytro (NDGF), Di (TRIUMF), Dave M (FNAL), Xin (BNL)

Experiments round table:

  • CMS reports ( raw view) -
    • Likely not able to attend the call
    • RAL FTS3 service in trouble on 25th(evening) until 26th (late morning) - GGUS:134744
      • FTS server not able to connect to various endpoints - no transfers on those links

  • ALICE -
    • Low activity Fri-Sat, else NTR

  • LHCb reports ( raw view) -
    • Activity
      • HLT farm off for MC
    • Updates
      • NTR
    • Site Issues
      • CERN: Staging issues. Many Connection reset by peer on Castor (GGUS:134755), related to FTS proxy-renewal. Noticed from 25th, bump of failures also this morning.

Sites / Services round table:

  • ASGC: nc
  • BNL: GGUS interface work ongoing, otherwise NTR
  • CNAF: nc
  • EGI: nc
  • FNAL: NTR
  • IN2P3: NTR
  • JINR: nc
  • KISTI: NTR
  • KIT: ntr
  • NDGF:
    • Last week we had two outages to our storage due to a card failure at the central router and subsequent cables fiddling. The first one (on Tuesday) got noticed, the second one (on Wednesday) apparently didn't.
    • NTR for this week
  • NL-T1: NTR
  • NRC-KI: nc
  • OSG: nc
  • PIC: nc
  • RAL: nc
  • TRIUMF: NTR

  • CERN computing services: nc
  • CERN storage services: NTR
  • CERN databases: Multiple DBoD instances in Wigner hosting IaaS data were affected by the NFS kernel bug and not available on Friday evening (OTG:0043694)
  • GGUS:
    • Added after the meeting:
      • As of today also team tickets have the downtime check popup window for the notified site
  • Monitoring: NTR
  • MW Officer: nc
  • Networks: nc
  • Security: nc

AOB:

Edit | Attach | Watch | Print version | History: r10 < r9 < r8 < r7 < r6 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r10 - 2018-04-30 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback